OpenStudy (anonymous):

The multiplication property of equality could be used to solve which of the following equations? (x + 2)(x - 3) = 0 m + 7 = -12 b ÷ (-2) = 18 -3 + y = 7 @koikkara

OpenStudy (koikkara):

@loveeemeee Hello, dear friend, Please ref: http://openstudy.com/study#/updates/50ac3e51e4b06b5e4933ac28

OpenStudy (anonymous):

geometry

OpenStudy (jchick):

The multiplication property of equality could be used to solve b ÷ (-2) = 18. b ÷ (-2) = 18 b ÷ (-2) * (-2) = 18 * (-2) b = -36